Using guass theorem calculate the flux of the vector field A =x³i+x²zj+yzk through the square of a cube of a side 2 unit's


1
Expert's answer
2021-05-13T09:21:08-0400

"\text{div} \vec{A}=3x^2\vec{i}+y\vec{k},"

"\Phi=\int _0^2dx\int_0^2dy\int _0^2(3x^2+y)dz=2\int_0^2dx\int_0^2(3x^2+y)dy=2\int_0^2(6x^2+2)dx=2(2x^3+2x)|_0^2=40~Wb."